OpenStudy (anonymous):

If f(x) = x + 2, find f(3). 1 6 5 −1

Parth (parthkohli):

\( \color{Black}{\Rightarrow f(x) = x + 2 \Longrightarrow f(3) = 3 + 2 }\)

OpenStudy (anonymous):

5